using the inductive reasoning to predict the next line in the pattern
(1x9)-6=3
(21x9)-6=183
(2*21x9)-6=2883
----
(3*21*9)-6 = 561
----------------------------
Cheers,
Stan H.
------------------

Answer by stanbon(75887)

using the inductive reasoning to predict the next line in the pattern
(1x9)-6=3
(21x9)-6=183
(2*21x9)-6=372
(3*21*9)-6 = 561
----
Cheers,
Stan H.
------------